# From Large to Small Datasets: Size Generalization for Clustering Algorithm Selection

### 22 Feb 2024 v1

@@ 2 Preliminaries
Let $\mathcal{X}=\{x_{i}\}_{i=1}^{n}$ be a clustering instance, $\mathcal{C}=(C_{1},...,C_{k})$ be a $k$-clustering of $\mathcal{S}\subset\mathcal{X}$, and $\mathcal{G}=(G_{1},...,G_{k})$ be a $k$-clustering of $\mathcal{X}$. For every $x_{i}\in\mathcal{X}$, let $\tau:\mathcal{X}\to[k]$ be a mapping such that $\tau(x)=j$ if and only if $x\in G_{j}$. We define the cost of clustering $\mathcal{C}$ with respect to the ground truth $\mathcal{G}$ as
${\sf{cost}}_{\mathcal{G}}\left(\mathcal{C}\right)\stackrel{{\scriptstyle\textup{def}}}{{=}}\frac{1}{\left\lvert\mathcal{S}\right\rvert}\min_{\sigma\in\Sigma_{k}}\sum_{x\in\mathcal{S}}\mathds{1}{\left\{x\in C_{\sigma(\tau(i))}\,\land\,x\notin G_{\tau(i)}\right\}}.$
@@ 3 $k$-Means++ and $k$-Centers Clustering
In this section, we provide size generalization bounds for $k$-means++ and $k$-centers clustering. Given $x\in\mathcal{X}$ and $C:=\{c_{1},...,c_{k}\}\subset\mathbb{R}^{d}$, let $d_{\sf center}(x;C)\stackrel{{\scriptstyle\textup{def}}}{{=}}\min_{c_{i}\in C}d(x,c_{i})$. The clustering of $\mathcal{X}$ induced by $C$ is the partition $S_{1},...,S_{k}$ where $S_{i}=\{x:{i=\operatorname*{argmin}}\left\lVert x-c_{j}\right\rVert\}$ (with fixed but arbitrary tie-breaking). The $k$-means problem minimizes the objective $d_{\sf center}(x;C)^{2}$, for which $k$-means++
provides a multiplicative $\log k$-approximation [1]. The $k$-centers problem minimizes the objective function $\max_{x\in\mathcal{X}}d_{\sf center}(x;C)$, for which Gonzalez’s heuristic [1985] is a multiplicative 2-approximation.
@@
Input: Instance $\mathcal{X}=(x_{1},...,x_{mk})\subset\mathbb{R}^{d}$, $k,m\in{\mathbb{Z}}_{\geq 0}$, $f:\mathbb{R}\times\mathcal{P}(\mathcal{X})\to\mathbb{R}_{\geq 0}$ Set $C^{1}=\{c_{1}\}$ with $c_{1}\sim\text{Unif}(\mathcal{X})$ for $i=2,3,\dots,k$ do
@@
Theorem 3.1 specifies the choice of $m$ in ${\sf ApxSeeding}$ that guarantees an $\epsilon$-approximation to the accuracy of ${\sf Seeding}$. Our bound depends linearly on $\zeta_{k,f}(\mathcal{X})$, a parameter [^2] To our knowledge, the parameter $\zeta_{k,f}(\mathcal{X})$ does not appear in previous literature, e.g., on coresets. However, $\zeta_{k,f}(\mathcal{X})$ can be viewed as a generalization of the parameter $\gamma^{\prime}$ from Theorem 1 by Bachem et al. [4]. measuring the smoothness of the distribution over centers induced by sampling according to $f$ in ${\sf Seeding}$ (line 1). Ideally, if this distribution induced by $f$ is close to uniform then $\zeta_{k,f}(\mathcal{X})$ is close to 1, but if it is very biased towards selecting certain points, then $\zeta_{k,f}(\mathcal{X})$ may be as large as $n$.
@@ 3.1 Application to $k$-Centers Clustering
where $f_{{\sf SoftmaxCenters}}(z;\mathcal{X})\stackrel{{\scriptstyle\textup{def}}}{{=}}\exp(\beta z)$. Theorem 3.2 shows that ${\sf SoftmaxCenters}$ provides a constant-factor approximation to the $k$-centers objective on instances with well-balanced $k$-centers solutions  [31, 22, 5]:
given $\mu_{\ell},\mu_{u}\in(0,1)$, we say a clustering instance $\mathcal{X}$ is ($\mu_{\ell},\mu_{u}$)-well-balanced with respect to a partition $S_{1},\dots,S_{k}$ if for all $i\in[k]$, $\mu_{\ell}|\mathcal{X}|\leq|{S}_{i}|\leq\mu_{u}|\mathcal{X}|$.
@@
$\zeta_{k,{\sf SoftmaxCenters}}(\mathcal{X})\stackrel{{\scriptstyle\textup{def}}}{{=}}n\cdot\max_{Q\subset\mathcal{X}:|Q|\leq k}\max_{x\in\mathcal{X}}\frac{\exp(\beta d_{\sf center}(x;Q))}{\sum_{y\in Q}\exp(\beta d_{\sf center}(x;Q))}\leq\exp(2\beta R).$
@@ 3.2 Application to $k$-Means++
Let $F$ be a distribution over $\mathbb{R}^{d}$ with variance $\sigma^{2}$ and support almost-surely bounded by an $R$-radius sphere. Suppose there is a $d^{\prime}$-dimensional, $r$-radius sphere $B$ such that $d^{\prime}>2$, $F(B)>0$, and $\forall x,y\in B$, $F(x)<cF(y)$ for some $c\geq 1$ (i.e., $F$ is non-degenerate). For $n$ sufficiently large, with high probability over $\mathcal{X}\sim F^{n}$,
@@ 4 Single Linkage Clustering
Input: Clustering instance $\mathcal{X}\subset\mathbb{R}^{d}$ and $k\in{\mathbb{Z}}_{\geq 0}$ Initialize $\mathcal{C}^{0}:=\{C^{0}_{1}=\{x_{1}\},...,C^{0}_{n}=\{x_{n}\}\}$, $i=1$ while $\left\lvert\mathcal{C}^{i}\right\rvert>1$ do
@@
             Choose $A,B\in\mathcal{C}^{i}$ such that $d(A,B)=d_{i}$ and set $\mathcal{C}^{i}=\mathcal{C}^{i}\setminus{\{A,B\}}\cup(A\cup B)$
@@
Our main result in this section bounds the number of samples sufficient and necessary to ensure that ${\sf SingleLinkage}(\mathcal{X}_{m},k)$ and ${\sf SingleLinkage}(\mathcal{X},k)$ return similar clusterings. Our bound depends on a parameter $\zeta_{k,{\sf SL}}\in(0,n]$, defined as follows.
with $\{C_{1},\dots,C_{k}\}={\sf SingleLinkage}(\mathcal{X},k)$:
@@
We emphasize that $\zeta_{k,{\sf SL}}$ measures the stability of ${\sf SingleLinkage}$ on $\mathcal{X}$ with respect to point deletion, but it is fundamentally unrelated to the cost of ${\sf SingleLinkage}$ on $\mathcal{X}$ with respect to a ground truth (which is what we ultimately aim to approximate). Any assumption that $\zeta_{k,{\sf SL}}$ is small relative to $n$ does not constitute any assumptions that ${\sf SingleLinkage}$ achieves good accuracy on $\mathcal{X}$.
Theorem 4.3 states our main result for size generalization of single linkage clustering.
@@
$m={\tilde{O}\left(\left(\frac{k}{\epsilon^{2}}+\frac{n}{\min_{i\in[k]}\left\lvert C_{i}\right\rvert^{2}}+\zeta_{{k,{\sf SL}}}(\mathcal{X})\right)\log\frac{k}{\delta}\right)},$
@@ C.1 Omitted Proofs from Section 3
$\hat{c}_{\mathcal{G}}(S^{\prime})\stackrel{{\scriptstyle\textup{def}}}{{=}}\min_{\sigma\in\Sigma_{k}}\frac{1}{\ell}\sum_{x\in\mathcal{X}^{\prime}}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(\tau(x))}\wedge x\in G_{\tau(x)}\right\}}.$
@@
$\operatorname*{\mathbb{E}}\left[\frac{1}{\ell}\sum_{x\in\mathcal{X}^{\prime}}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(\tau(x))}\wedge x\in G_{\tau(x)}\right\}}\right]=\frac{1}{n}\sum_{x\in\mathcal{X}}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(\tau(x))}\wedge x\in G_{\tau(i)}\right\}}.$
@@
$\frac{1}{\ell}\sum_{x\in\mathcal{X}^{\prime}}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(\tau(x))}\wedge x\in G_{\tau(x)}\right\}}\approx_{\epsilon}\frac{1}{n}\sum_{x\in\mathcal{X}}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(\tau(x))}\wedge x\in G_{\tau(i)}\right\}}$
@@ C.2 Omitted Proofs from Section 4
$\frac{1}{m}\sum_{x\in\mathcal{X}_{m}}\mathds{1}{\left\{x\in C^{\prime}_{\sigma(\tau(x))}\wedge x\notin G_{\tau(x)}\right\}}$ $=\frac{1}{m}\sum_{x\in\mathcal{X}}\mathds{1}{\left\{x\in C^{\prime}_{\sigma(\tau(x))}\wedge x\notin G_{\tau(x)}\wedge x\in\mathcal{X}_{m}\right\}}$
$=\frac{1}{m}\sum_{x\in\mathcal{X}}\mathds{1}{\left\{x\in C_{\sigma(\tau(x))}\wedge x\notin G_{\tau(x)}\wedge x\in\mathcal{X}_{m}\right\}},$
$\operatorname*{\mathbb{E}}\left[\frac{1}{m}\sum_{x\in\mathcal{X}_{m}}\mathds{1}{\left\{x\in C^{\prime}_{\sigma(\tau(x))}\wedge x\notin G_{\tau(x)}\right\}}\right]$ $=\operatorname*{\mathbb{E}}\left[\frac{1}{m}\sum_{x\in\mathcal{X}}\mathds{1}{\left\{x\in C_{\sigma(\tau(x))}\wedge x\notin G_{\tau(x)}\wedge x\in\mathcal{X}_{m}\right\}}\right]$
$=\frac{1}{m}\frac{m}{n}\cdot\operatorname*{\mathbb{E}}\left[\sum_{x\in\mathcal{X}}\mathds{1}{\left\{x\in C_{\sigma(\tau(x))}\wedge x\in G_{\tau(x)}\right\}}\right].$

### 26 Feb 2024 v2

@@ 2 Preliminaries
Let $\mathcal{X}=\{x_{i}\}_{i=1}^{n}$ be a clustering instance, $\mathcal{C}=(C_{1},...,C_{k})$ be a $k$-clustering of $\mathcal{S}\subset\mathcal{X}$, and $\mathcal{G}=(G_{1},...,G_{k})$ be a $k$-clustering of $\mathcal{X}$. We define the cost of clustering $\mathcal{C}$ with respect to the ground truth $\mathcal{G}$ as
${\sf{cost}}_{\mathcal{G}}\left(\mathcal{C}\right)\stackrel{{\scriptstyle\textup{def}}}{{=}}\frac{1}{\left\lvert\mathcal{S}\right\rvert}\min_{\sigma\in\Sigma_{k}}\sum_{x\in\mathcal{S}}\sum_{j\in[k]}\mathds{1}{\left\{x\in C_{\sigma(j)}\,\land\,x\notin G_{j}\right\}}.$
@@ 3 $k$-Means++ and $k$-Centers Clustering
In this section, we provide size generalization bounds for $k$-means++ and $k$-centers clustering. Given $x\in\mathcal{X}$ and $C:=\{c_{1},...,c_{k}\}\subset\mathbb{R}^{d}$, let $d_{\sf center}(x;C)\stackrel{{\scriptstyle\textup{def}}}{{=}}\min_{c_{i}\in C}d(x,c_{i})$. The clustering of $\mathcal{X}$ induced by $C$ is the partition $S_{1},...,S_{k}$ where $S_{i}=\{x:{i=\operatorname*{argmin}}\left\lVert x-c_{j}\right\rVert\}$ (with fixed but arbitrary tie-breaking). The $k$-means problem minimizes the objective $\sum_{x\in\mathcal{X}}d_{\sf center}(x;C)^{2}$, for which $k$-means++ provides a multiplicative $\log k$-approximation [1]. The $k$-centers problem minimizes the objective function $\max_{x\in\mathcal{X}}d_{\sf center}(x;C)$, for which Gonzalez’s heuristic [1985] is a multiplicative 2-approximation.
@@
Input: Instance $\mathcal{X}=(x_{1},...,x_{mk})\subset\mathbb{R}^{d}$, $k,m\in{\mathbb{Z}}_{\geq 0}$, $f:\mathbb{R}\times\mathcal{P}(\mathcal{X})\to\mathbb{R}_{\geq 0}$ Set $C^{1}=\{c_{1}\}$ with $c_{1}\sim\text{Unif}(\mathcal{X})$ $\ell=1$ for $i=2,3,\dots,k$ do
@@
Theorem 3.1 specifies the choice of $m$ in ${\sf ApxSeeding}$ that guarantees an $\epsilon$-approximation to the accuracy of ${\sf Seeding}$. Our bound depends linearly on $\zeta_{k,f}(\mathcal{X})$, a parameter [^2] To our knowledge, the parameter $\zeta_{k,f}(\mathcal{X})$ does not appear in previous literature, e.g., on coresets. However, $\zeta_{k,f}(\mathcal{X})$ can be viewed as a generalization of the parameter $\gamma^{\prime}$ from Theorem 1 by Bachem et al. [4]. measuring the smoothness of the distribution over centers induced by sampling according to $f$ in ${\sf Seeding}$ (line 1). Ideally, if this distribution induced by $f$ is close to uniform, then $\zeta_{k,f}(\mathcal{X})$ is close to 1, but if it is very biased towards selecting certain points, then $\zeta_{k,f}(\mathcal{X})$ may be as large as $n$.
@@ 3.1 Application to $k$-Centers Clustering
where $f_{{\sf SoftmaxCenters}}(z;\mathcal{X})\stackrel{{\scriptstyle\textup{def}}}{{=}}\exp(\beta z)$. Theorem 3.2 shows that ${\sf SoftmaxCenters}$ provides a constant-factor approximation to the $k$-centers objective on instances with well-balanced $k$-centers solutions  [31, 22, 5]: given $\mu_{\ell},\mu_{u}\in(0,1)$, we say a clustering instance $\mathcal{X}$ is ($\mu_{\ell},\mu_{u}$)-well-balanced with respect to a partition $S_{1},\dots,S_{k}$ if for all $i\in[k]$, $\mu_{\ell}|\mathcal{X}|\leq|{S}_{i}|\leq\mu_{u}|\mathcal{X}|$.
@@
$\zeta_{k,{\sf SoftmaxCenters}}(\mathcal{X})\stackrel{{\scriptstyle\textup{def}}}{{=}}n\cdot\max_{Q\subset\mathcal{X}:|Q|\leq k}\max_{x\in\mathcal{X}}\frac{\exp(\beta d_{\sf center}(x;Q))}{\sum_{y\in Q}\exp(\beta d_{\sf center}(y;Q))}\leq\exp(2\beta R).$
@@ 3.2 Application to $k$-Means++
Let $F$ be a distribution over $\mathbb{R}^{d}$ with expectation $\tilde{\mu}\in\mathbb{R}^{d}$ and variance $\sigma^{2}\stackrel{{\scriptstyle\textup{def}}}{{=}}\int_{x}d(x,\tilde{\mu})^{2}F(x)$ and support almost-surely bounded by an $R$-radius sphere. Suppose there is a $d^{\prime}$-dimensional, $r$-radius sphere $B$ such that $d^{\prime}>2$, $F(B)>0$, and $\forall x,y\in B$, $F(x)<cF(y)$ for some $c\geq 1$ (i.e., $F$ is non-degenerate). For $n$ sufficiently large, with high probability over $\mathcal{X}\sim F^{n}$,
@@ 4 Single Linkage Clustering
Input: Clustering instance $\mathcal{X}\subset\mathbb{R}^{d}$ and $k\in{\mathbb{Z}}_{\geq 0}$ Initialize $\mathcal{C}^{0}:=\{C^{0}_{1}=\{x_{1}\},...,C^{0}_{n}=\{x_{n}\}\}$, $i=1$ while $\left\lvert\mathcal{C}^{i-1}\right\rvert>1$  do
@@
             Choose $A,B\in\mathcal{C}^{i}$ such that $d(A,B)=d_{i}$ and set $\mathcal{C}^{i}=\mathcal{C}^{i}\setminus{\{A,B\}}\cup\{A\cup B\}$
@@
Our main result in this section bounds the number of samples sufficient and necessary to ensure that ${\sf SingleLinkage}(\mathcal{X}_{m},k)$ and ${\sf SingleLinkage}(\mathcal{X},k)$ return similar clusterings. Our bound depends on a parameter $\zeta_{k,{\sf SL}}\in(0,n]$, defined as follows, with $\{C_{1},\dots,C_{k}\}={\sf SingleLinkage}(\mathcal{X},k)$:
@@
We emphasize that $\zeta_{k,{\sf SL}}$ measures the stability of ${\sf SingleLinkage}$ on $\mathcal{X}$ with respect to point deletion, but it is fundamentally unrelated to the cost of ${\sf SingleLinkage}$ on $\mathcal{X}$ with respect to a ground truth (which is what we ultimately aim to approximate). Any assumption that $\zeta_{k,{\sf SL}}$ is small relative to $n$ does not constitute any assumptions that ${\sf SingleLinkage}$ achieves good accuracy on $\mathcal{X}$. Theorem 4.3 states our main result for size generalization of single linkage clustering.
@@
$m={\tilde{O}\left(\left(\frac{k}{\epsilon^{2}}+\frac{n}{\min_{i\in[k]}\left\lvert C_{i}\right\rvert}+\zeta_{{k,{\sf SL}}}(\mathcal{X})\right)\log\frac{k}{\delta}\right)},$
@@ C.1 Omitted Proofs from Section 3
$\hat{c}_{\mathcal{G}}(S^{\prime})\stackrel{{\scriptstyle\textup{def}}}{{=}}\min_{\sigma\in\Sigma_{k}}\frac{1}{\ell}\sum_{x\in\mathcal{X}^{\prime}}\sum_{i\in[k]}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\right\}}.$
@@
$\operatorname*{\mathbb{E}}\left[\frac{1}{\ell}\sum_{x\in\mathcal{X}^{\prime}}\sum_{i\in[k]}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\right\}}\right]=\frac{1}{n}\sum_{x\in\mathcal{X}}\sum_{i\in[k]}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\right\}}.$
@@
$\frac{1}{\ell}\sum_{x\in\mathcal{X}^{\prime}}\sum_{i\in[k]}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\right\}}\approx_{\epsilon}\frac{1}{n}\sum_{x\in\mathcal{X}}\sum_{i\in[k]}\mathds{1}{\left\{x\in S^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\right\}}$
@@ C.2 Omitted Proofs from Section 4
$\frac{1}{m}\sum_{x\in\mathcal{X}_{m}}\sum_{i\in[k]}\mathds{1}{\left\{x\in C^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\right\}}$ $=\frac{1}{m}\sum_{x\in\mathcal{X}}\sum_{i\in[k]}\mathds{1}{\left\{x\in C^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\wedge x\in\mathcal{X}_{m}\right\}}$
$=\frac{1}{m}\sum_{x\in\mathcal{X}}\sum_{i\in[k]}\mathds{1}{\left\{x\in C_{\sigma(i)}\wedge x\notin G_{i}\wedge x\in\mathcal{X}_{m}\right\}},$
$\operatorname*{\mathbb{E}}\left[\frac{1}{m}\sum_{x\in\mathcal{X}_{m}}\sum_{i\in[k]}\mathds{1}{\left\{x\in C^{\prime}_{\sigma(i)}\wedge x\notin G_{i}\right\}}\right]$ $=\operatorname*{\mathbb{E}}\left[\frac{1}{m}\sum_{x\in\mathcal{X}}\sum_{i\in[k]}\mathds{1}{\left\{x\in C_{\sigma(i)}\wedge x\notin G_{i}\wedge x\in\mathcal{X}_{m}\right\}}\right]$
$=\frac{1}{m}\frac{m}{n}\cdot\operatorname*{\mathbb{E}}\left[\sum_{x\in\mathcal{X}}\sum_{i\in[k]}\mathds{1}{\left\{x\in C_{\sigma(i)}\wedge x\notin G_{i}\right\}}\right].$